Is f(x)=1/x2/3 in L3[1 2] ?


1
Expert's answer
2022-01-12T17:18:20-0500

for L3[1 2] :

"\int^2_1|f(x)|^3dx<\infin"


then we have:

"\int^2_1x^{-2}dx=-1/x|^2_1=-1/2+1=1/2<\infin"


so,  f(x)=1/x2/3 in L3[1 2]